# Near-Inertial Pollard Waves Modeling the Arctic Halocline

### v3

@@
Abstract.
We present an explicit and exact solution to the governing equations describing the vertical structure of the Arctic Ocean region centred around the North Pole. The solution describes a stratified water column with three constant-density regions: a motionless bottom layer, a middle layer - the halocline - described by nonhydrostatic, near-inertial Pollard waves, and an upper layer presenting a mean current and a wave motion associated with the one in the halocline layer.
Key words and phrases:
Euler equations; Halocline; Arctic Ocean; North Pole; Pollard waves; Gerstner waves
1991 Mathematics Subject Classification:
35Q35; 76U60; 35C05; 35Q31; 37N10
ORCID: 0009-0008-5454-0922
@@ 1. Introduction
The Arctic Ocean, located in the Northern Hemisphere and encompassing the North Pole, is a small ocean about $4000\,\mathrm{km}$ long and $2500\,\mathrm{km}$ wide, covering an area of around 10 million $\mathrm{km}^{2}$. It includes two main deep basins (around $4000\,\mathrm{m}$ deep) separated by the underwater Lomonosov Ridge, the Amerasian Basin (divided into the Makarov and Canada basins), and the smaller Eurasian Basin (divided into the Amundsen and Nansen basins), and is surrounded by shallow seas (less than $400\,\mathrm{m}$ deep): the Barents Sea, the Kara Sea, the Laptev Sea, the Siberian Sea, the Chukchi Sea and the Beaufort Sea (see [BertosioPHD] or [Talley]). See Figure 1.1.
@@
Figure 1.1. Map of the Arctic Ocean. Adapted from the IBCAO map https://www.ngdc.noaa.gov/mgg/bathymetry/arctic/currentmap.html. Credit [IBCAO]
@@
The primary oceanic inflows are from the Atlantic Ocean through the Fram Strait and the Barents Sea, and from the Pacific Ocean via the Bering Strait. Significant freshwater inflows also come from rivers in North America and Siberia. These inflows are major factors influencing the salinity and temperature variability of the water column, with salinity playing a key role in stratification, creating a halocline rather than a thermocline (which is instead typical of mid-latitudes and equatorial regions). As a matter of fact, one of the main features of the Arctic Ocean is that it is predominantly stratified by salinity (hence it is defined as a $\beta$-ocean) rather than by temperature (the so called $\alpha$-oceans) (see[BertosioPHD] or [Carmack2007]).
@@
Moreover, the central region of Arctic Ocean - centred around the North Pole and subject of our analysis - is entirely covered by a relatively thin (not exceeding $2\,\mathrm{m}$ of thickness) layer of sea ice, which reduces in summer when temperature increases. Due to the global warming, the sea ice is rapidly declining, and the Arctic region is warming faster than the global average, due to the process known as Arctic amplification. This makes the Arctic highly vulnerable to climate change, potentially leading to an ice-free Arctic Ocean in summer and with thinner and more mobile sea ice in winter. The transition to a seasonally ice-free Arctic will profoundly affect Arctic oceanography, the marine ecosystems it supports, and the global climate (see [TM2020]).
@@
In addition to the atmospheric conditions, the Arctic sea ice is influenced by the sea water beneath it. The Arctic Ocean is stratified into a cold and fresh surface mixed layer (SML), with a depth between $5$ to $100$ meters, a halocline below the mixed layer with a base depth ranging from $40$ to $200$ meters, and a layer of warmer and saltier Atlantic Water (AW) (see [MetznerSalzmann]). The depths of the boundaries of this layers varies mainly according to temperature and presence of ice. In general, it can be said that in winter, with lower temperatures and increased ice presence, their depth increases, whereas in summer, these decrease (see [Peralta]).
@@
The halocline is a region of strong stratification, which prevents interaction of the ice cover with AW heat by the direct surface-generated mixing of the SML (see [Polyakov2018] and [Guthrie] for an in-detail description of the Arctic water column and its main physical processes); consequently, the halocline layer, situated above the saltier and denser water, is fundamental for the formation of the ice cover (see [Polyakov2018]).
@@
The physical processes in the mixed layer, mainly the ice-motion induced by wind and the Transpolar Drift Current (TDC) - the Arctic Ocean current transporting surface waters and sea ice from the Laptev Sea and the East Siberian Sea towards Fram Strait, with a speed of around $0.07\,\mathrm{ms^{-1}}$ around the North Pole - induces a current with magnitude approximately $0.1\,\mathrm{ms^{-1}}$ on the lower boundary of the mixed layer (see [Guthrie] and data therein).
@@
The solution in this study is constructed by adapting Pollard’s solution, presented in [Pollard] for surface waves accounting for the effects of Earth’s rotation and extending the remarkable solution provided by Gerstner in [Gerstner]. See also [C2001], [ConstantinBook] and [Henry], and [CW] for more general trochoidal solutions. Recently, Gerstner-like solutions were used to describe equatorial waves [Constantin2012], [Constantin2013], [Constantin2014], [ConstatinEquatorial], [Henry16] and to study their linear instability [CG], [HStability], [HC15], or to study wave-current interactions [Constantin2017], [K17] [McCarney2023], [McCarney2024] and edge waves [CEdgewaves], [IK2015] [Weber], [Miao]. See also [Hsu14], [R-S2018], [Hsu17] and [ChuETAL] for Gerstner-like waves in other oceanographic models.
Similarly, Pollard-like internal wave approaches have been developed in [IK15] or [K19], and their instability has also been subject of investigation [IKshortw], [IKstability].
Despite the fact that within the halocline one could distinguish between the cold halocline layer in the Eurasian Basin, the Pacific Halocline Waters in the Amerasian Basin and the lower halocline water (see [MetznerSalzmann]), the simplified model we will consider features constant densities in the layers under consideration. The case of depth-dependent densities could be considered in future works, even if density variations are very small within each stratum (see [Polyakov2018]), so accounting for density variation would probably not provide substantial differences. Our model features three constant densities, $\rho_{0}$, $\rho_{1}$ and $\rho_{2}$, with $\rho_{0}<\rho_{1}<\rho_{2}$, where $\rho_{0}$ represents the fresh and cold water of the surface mixed layer above the halocline, $\rho_{1}$ is the density of the halocline’s water and $\rho_{2}$ is the density of the saltier and warmer water (AW) below the halocline. See Figure 1.2.
Figure 1.2. A depiction of the model we are considering
More precisely, we assume the following water characteristics (we refer to [Rudels]):
@@
and density variations can be modeled by (see [BertosioPHD] or [Talley])
@@
where $\mathfrak{T}$ is the potential temperature, $\mathfrak{S}$ is the salinity, $\alpha\approx 53\cdot 10^{-6}\,\mathrm{K^{-1}}$ is the thermal expansion coefficient and $\beta\approx 785\cdot 10^{-6}\,\mathrm{kg}\,\mathrm{g^{-1}}$ is the haline contraction coefficient. Both $\alpha$ and $\beta$ depend of the water properties, therefore the value reported here refer only for the Arctic Ocean (see [Talley]).
@@
  • in Section 2 we derive the governing equations. More precisely, as we are investigating the fluid dynamics around the North Pole, where the classical spherical coordinates fail, we need to adopt the rotated spherical coordinates developed in [CJ2023], and therefore adapt the tangent-plane, traditional and $f$-plane approximations to this new coordinate system, starting from the Euler equations. Then, by a $2D$ rotation we will align one of the axes with the Transpolar Drift Current in order to simplify the study of the internal waves propagating in its direction;
@@
    In previous works [Constantin2017], [Kluczek], [McCarney2024], with only one dynamic boundary condition, two modes of the wave motion were found, one fast mode standard to the theory of internal waves, and a second one, slow, with the period close to the inertial period of the Earth $T_{i}=\frac{2\pi}{f}$. Constantin & Monismith [Constantin2017] refer to this slow-mode wave as inertial Gerstner wave.
@@ 2.1. Classical and rotated spherical coordinates
We consider the Earth to be a sphere of radius $R\approx 6371\,\mathrm{km}$. The classical spherical coordinate system is not suitable for the study of flows in regions centred around the North (and South) Pole, as longitude is undefined at the poles. This result is a consequence of the ‘hairy ball theorem’ (see [Chinn] and [CJBook]). In this section, we review the construction of the rotated spherical coordinate system developed in [CJ2023] (see also [CJBook] for a detailed exposition) to avoid this problem, and we derive the $f$-plane approximation for the Euler equations in this new coordinate system.
@@
and the corresponding velocity components are $\boldsymbol{u}=(u^{{\dagger}},v^{{\dagger}},w^{{\dagger}})$. The two coordinate systems are depicted in Figure 2.1.
Figure 2.1. Left: classical spherical coordinates. Right: rotated spherical coordinate. Images from [CJ2023]. $C$ BY 4.0 (http://creativecommons.org/licenses/by/4.0)
@@ 2.2. Equations of motion
Away from the boundary layers, the friction/viscous effects are negligible, so the fluid can be considered ideal [CJBOOK]. The governing equations for a rotating incompressible ideal fluid are, in vectorial coordinate-free form, written as
@@
where $\frac{D}{Dt}=\left(\frac{\partial}{\partial t}+\boldsymbol{u}\cdot\nabla\right)$ is the material derivative, $\boldsymbol{u}$ is the velocity vector, $\rho$ is the fluid density, $\mathbf{G}$ is the body force per unit volume, $p$ is the pressure, $\mathbf{\Omega}$ the rotation vector and $\mathbf{r}$ the position vector [ConstantinBook], [Morita]. In the subsequent analysis, we will consider three constant densities $\rho_{0}$, $\rho_{1}$ and $\rho_{2}$, with $\rho_{0}<\rho_{1}<\rho_{2}$, in different regions of the flow; in each layer the fluid can be considered incompressible.
@@
In the new rotated spherical coordinates the Euler equations are written as (see [MIO] for a complete derivation of the Navier-Stokes equations in the rotated spherical coordinate system, from which the incompressible Euler equations reduces to, when considering constant density and zero viscosity)
@@ 2.2.1. Standard approximations in geophysical fluid dynamics
In this paragraph, we review the ideas leading to the so-called traditional approximation, adapting them to the rotated spherical coordinates. Namely, we will construct this approximation starting from (2.13). Note that, even if the velocities $u^{{\dagger}},v^{{\dagger}},w^{{\dagger}}$ are defined differently from the ones relative to the classical spherical coordinate system, $u^{{\dagger}},v^{{\dagger}}$ are still the horizontal and $w^{{\dagger}}$ the vertical components of the velocity vector $\boldsymbol{u}$, hence the observations about their magnitude, on which the traditional approximation is based, still apply (see [Vallis]). For a review of the traditional approximation of the governing equations in classical spherical coordinates, see also [MarshallPlumb].
@@
These two approximations related to the Coriolis term $2\mathbf{\Omega}\times\mathbf{u}$, the first involving the two horizontal components of the equation, and the second involving the vertical component where the gravity acceleration is present, cannot be done independently of each other. Instead, they must be made together, otherwise the resulting equations fail to conserve energy [Salmon].
Lastly, as the ocean is shallow compared to Earth’s radius, we can write $r=R+\mathsf{z}$, with $R$ being the Earth’s radius and $\mathsf{z}$ increasing in the radial direction. This idea is fundamental for the so-called ‘thin-shell’ approximation (see e.g [CJ2023] or [JBook]). The coordinate $r$ is then replaced by $R$, except in the derivatives. For example $\frac{1}{r^{2}}\frac{\partial r^{2}w^{{\dagger}}}{\partial r}$ becomes $\frac{\partial w^{{\dagger}}}{\partial\mathsf{z}}$. With the above arguments, the Euler, mass conservation and incompressibility equations reduce to
$\left\{\begin{aligned} &\left[\frac{\partial}{\partial t}+\frac{u^{{\dagger}}}{R\cos\theta^{{\dagger}}}\frac{\partial}{\partial\varphi^{{\dagger}}}+\frac{v^{{\dagger}}}{R}\frac{\partial}{\partial\theta^{{\dagger}}}+w^{{\dagger}}\frac{\partial}{\partial\mathsf{z}}\right]\begin{pmatrix}u^{{\dagger}}\\ v^{{\dagger}}\\ w^{{\dagger}}\\ \end{pmatrix}+\\ &\qquad+2\Omega\begin{pmatrix}-v^{{\dagger}}\sin\varphi^{{\dagger}}\cos\theta^{{\dagger}}\\ u^{{\dagger}}\sin\varphi^{{\dagger}}\cos\theta^{{\dagger}}\\ 0\end{pmatrix}=-\frac{1}{\rho}\begin{pmatrix}\frac{1}{R\cos\theta^{{\dagger}}}\frac{\partial P}{\partial\varphi^{{\dagger}}}\\ \frac{1}{R}\frac{\partial P}{\partial\theta^{{\dagger}}}\\ \frac{\partial P}{\partial\mathsf{z}}\\ \end{pmatrix}-\begin{pmatrix}0\\ 0\\ g\\ \end{pmatrix}\\ &\,\frac{\partial\rho}{\partial t}+\frac{u^{{\dagger}}}{R\cos\theta^{{\dagger}}}\frac{\partial\rho}{\partial\varphi^{{\dagger}}}+\frac{v^{{\dagger}}}{R}\frac{\partial\rho}{\partial\theta^{{\dagger}}}+w^{{\dagger}}\frac{\partial\rho}{\partial\mathsf{z}}=0\\ &\,\frac{1}{R\cos\theta^{{\dagger}}}\frac{\partial u^{{\dagger}}}{\partial\varphi^{{\dagger}}}+\frac{1}{R}\frac{\partial v^{{\dagger}}}{\partial\theta^{{\dagger}}}+\frac{\partial w^{{\dagger}}}{\partial\mathsf{z}}=0,\end{aligned}\right$. (2.23)
@@
Usually, the Coriolis parameter is defined by $f=2\Omega\sin\theta$, where $\theta\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ is the classical angle of latitude (see [Morita]), but, due to the relation $\sin\theta=\cos\theta^{{\dagger}}\sin\varphi^{{\dagger}}$ in (2.5) we can write

### v4

@@
Abstract. We present an explicit and exact solution to the governing equations describing the vertical structure of the Arctic Ocean region centred around the North Pole. The solution describes a stratified water column with three constant-density regions: a motionless bottom layer, a middle layer—the halocline—described by nonhydrostatic, near-inertial Pollard waves, and an upper layer presenting a mean current and a wave motion associated with the one in the halocline layer.
@@ 1. Introduction
The Arctic Ocean, located in the Northern Hemisphere and encompassing the North Pole, is a small ocean about $4000\,\mathrm{km}$ long and $2500\,\mathrm{km}$ wide, covering an area of around 10 million $\mathrm{km}^{2}$. It includes two main deep basins (around $4000\,\mathrm{m}$ deep) separated by the underwater Lomonosov Ridge, the Amerasian Basin (divided into the Makarov and Canada basins), and the smaller Eurasian Basin (divided into the Amundsen and Nansen basins), and is surrounded by shallow seas (less than $400\,\mathrm{m}$ deep): the Barents Sea, the Kara Sea, the Laptev Sea, the Siberian Sea, the Chukchi Sea and the Beaufort Sea (see [2] or [55]). See Figure 1.
@@
Figure 1. Map of the Arctic Ocean. Adapted from the IBCAO map https://www.ngdc.noaa.gov/mgg/bathymetry/arctic/currentmap.html. Credit [33].
@@
The primary oceanic inflows are from the Atlantic Ocean through the Fram Strait and the Barents Sea, and from the Pacific Ocean via the Bering Strait. Significant freshwater inflows also come from rivers in North America and Siberia. These inflows are major factors influencing the salinity and temperature variability of the water column, with salinity playing a key role in stratification, creating a halocline rather than a thermocline (which is instead typical of mid-latitudes and equatorial regions). As a matter of fact, one of the main features of the Arctic Ocean is that it is predominantly stratified by salinity (hence it is defined as a $\beta$-ocean) rather than by temperature (the so called $\alpha$-oceans) (see[2] or [3]).
@@
Moreover, the central region of Arctic Ocean—centred around the North Pole and subject of our analysis—is entirely covered by a relatively thin (not exceeding $2\,\mathrm{m}$ of thickness) layer of sea ice, which reduces in summer when temperature increases. Due to the global warming, the sea ice is rapidly declining, and the Arctic region is warming faster than the global average, due to the process known as Arctic amplification. This makes the Arctic highly vulnerable to climate change, potentially leading to an ice-free Arctic Ocean in summer and with thinner and more mobile sea ice in winter. The transition to a seasonally ice-free Arctic will profoundly affect Arctic oceanography, the marine ecosystems it supports, and the global climate (see [56]).
@@
In addition to the atmospheric conditions, the Arctic sea ice is influenced by the sea water beneath it. The Arctic Ocean is stratified into a cold and fresh surface mixed layer (SML), with a depth between $5$ to $100$ meters, a halocline below the mixed layer with a base depth ranging from $40$ to $200$ meters, and a layer of warmer and saltier Atlantic Water (AW) (see [41]). The depths of the boundaries of this layers varies mainly according to temperature and presence of ice. In general, it can be said that in winter, with lower temperatures and increased ice presence, their depth increases, whereas in summer, these decrease (see [46]).
@@
The halocline is a region of strong stratification, which prevents interaction of the ice cover with AW heat by the direct surface-generated mixing of the SML (see [49] and [22] for an in-detail description of the Arctic water column and its main physical processes); consequently, the halocline layer, situated above the saltier and denser water, is fundamental for the formation of the ice cover (see [49]).
@@
The physical processes in the mixed layer, mainly the ice-motion induced by wind and the Transpolar Drift Current (TDC)—the Arctic Ocean current transporting surface waters and sea ice from the Laptev Sea and the East Siberian Sea towards Fram Strait, with a speed of around $0.07\,\mathrm{ms^{-1}}$ around the North Pole—induces a current with magnitude approximately $0.1\,\mathrm{ms^{-1}}$ on the lower boundary of the mixed layer (see [22] and data therein).
@@
The solution in this study is constructed by adapting Pollard’s solution, presented in [48] for surface waves accounting for the effects of Earth’s rotation and extending the remarkable solution provided by Gerstner in [21]. See also [6], [8] and [23], and [18] for more general trochoidal solutions. Recently, Gerstner-like solutions were used to describe equatorial waves [10], [11], [12], [9], [24] and to study their linear instability [16], [26], [25], or to study wave-current interactions [17], [35] [42], [43] and edge waves [7], [30] [58], [44]. See also [27], [51], [28] and [5] for Gerstner-like waves in other oceanographic models.
Similarly, Pollard-like internal wave approaches have been developed in [29] or [38], and their instability has also been subject of investigation [32], [31].
Despite the fact that within the halocline one could distinguish between the cold halocline layer in the Eurasian Basin, the Pacific Halocline Waters in the Amerasian Basin and the lower halocline water (see [41]), the simplified model we will consider features constant densities in the layers under consideration. The case of depth-dependent densities could be considered in future works, even if density variations are very small within each stratum (see [49]), so accounting for density variation would probably not provide substantial differences. Our model features three constant densities, $\rho_{0}$, $\rho_{1}$ and $\rho_{2}$, with $\rho_{0}<\rho_{1}<\rho_{2}$, where $\rho_{0}$ represents the fresh and cold water of the surface mixed layer above the halocline, $\rho_{1}$ is the density of the halocline’s water and $\rho_{2}$ is the density of the saltier and warmer water (AW) below the halocline. See Figure 2.
Figure 2. A depiction of the model we are considering.
More precisely, we assume the following water characteristics (we refer to [52]):
@@
and density variations can be modeled by (see [2] or [55])
@@
where $\mathfrak{T}$ is the potential temperature, $\mathfrak{S}$ is the salinity, $\alpha\approx 53\cdot 10^{-6}\,\mathrm{K^{-1}}$ is the thermal expansion coefficient and $\beta\approx 785\cdot 10^{-6}\,\mathrm{kg}\,\mathrm{g^{-1}}$ is the haline contraction coefficient. Both $\alpha$ and $\beta$ depend of the water properties, therefore the value reported here refer only for the Arctic Ocean (see [55]).
@@
  • in Section 2 we derive the governing equations. More precisely, as we are investigating the fluid dynamics around the North Pole, where the classical spherical coordinates fail, we need to adopt the rotated spherical coordinates developed in [14], and therefore adapt the tangent-plane, traditional and $f$-plane approximations to this new coordinate system, starting from the Euler equations. Then, by a $2D$ rotation we will align one of the axes with the Transpolar Drift Current in order to simplify the study of the internal waves propagating in its direction;
@@
    In previous works [17], [37], [43], with only one dynamic boundary condition, two modes of the wave motion were found, one fast mode standard to the theory of internal waves, and a second one, slow, with the period close to the inertial period of the Earth $T_{i}=\frac{2\pi}{f}$. Constantin & Monismith [17] refer to this slow-mode wave as inertial Gerstner wave.
@@ 2.1. Classical and rotated spherical coordinates
We consider the Earth to be a sphere of radius $R\approx 6371\,\mathrm{km}$. The classical spherical coordinate system is not suitable for the study of flows in regions centred around the North (and South) Pole, as longitude is undefined at the poles. This result is a consequence of the ‘hairy ball theorem’ (see [4] and [15]). In this section, we review the construction of the rotated spherical coordinate system developed in [14] (see also [15] for a detailed exposition) to avoid this problem, and we derive the $f$-plane approximation for the Euler equations in this new coordinate system.
@@
and the corresponding velocity components are $\boldsymbol{u}=(u^{{\dagger}},v^{{\dagger}},w^{{\dagger}})$. The two coordinate systems are depicted in Figure 3.
Figure 3. Left: classical spherical coordinates. Right: rotated spherical coordinate. Images from [14]. $C$ BY 4.0 (http://creativecommons.org/licenses/by/4.0).
@@ 2.2. Equations of motion
Away from the boundary layers, the friction/viscous effects are negligible, so the fluid can be considered ideal [13]. The governing equations for a rotating incompressible ideal fluid are, in vectorial coordinate-free form, written as
@@
where $\frac{D}{Dt}=\left(\frac{\partial}{\partial t}+\boldsymbol{u}\cdot\nabla\right)$ is the material derivative, $\boldsymbol{u}$ is the velocity vector, $\rho$ is the fluid density, $\mathbf{G}$ is the body force per unit volume, $p$ is the pressure, $\mathbf{\Omega}$ the rotation vector and $\mathbf{r}$ the position vector [8], [45]. In the subsequent analysis, we will consider three constant densities $\rho_{0}$, $\rho_{1}$ and $\rho_{2}$, with $\rho_{0}<\rho_{1}<\rho_{2}$, in different regions of the flow; in each layer the fluid can be considered incompressible.
@@
In the new rotated spherical coordinates the Euler equations are written as (see [50] for a complete derivation of the Navier-Stokes equations in the rotated spherical coordinate system, from which the incompressible Euler equations reduces to, when considering constant density and zero viscosity)
@@ 2.2.1. Standard approximations in geophysical fluid dynamics
In this paragraph, we review the ideas leading to the so-called traditional approximation, adapting them to the rotated spherical coordinates. Namely, we will construct this approximation starting from (2.13). Note that, even if the velocities $u^{{\dagger}},v^{{\dagger}},w^{{\dagger}}$ are defined differently from the ones relative to the classical spherical coordinate system, $u^{{\dagger}},v^{{\dagger}}$ are still the horizontal and $w^{{\dagger}}$ the vertical components of the velocity vector $\boldsymbol{u}$, hence the observations about their magnitude, on which the traditional approximation is based, still apply (see [57]). For a review of the traditional approximation of the governing equations in classical spherical coordinates, see also [40].
@@
These two approximations related to the Coriolis term $2\mathbf{\Omega}\times\mathbf{u}$, the first involving the two horizontal components of the equation, and the second involving the vertical component where the gravity acceleration is present, cannot be done independently of each other. Instead, they must be made together, otherwise the resulting equations fail to conserve energy [53].
Lastly, as the ocean is shallow compared to Earth’s radius, we can write $r=R+\mathsf{z}$, with $R$ being the Earth’s radius and $\mathsf{z}$ increasing in the radial direction. This idea is fundamental for the so-called ‘thin-shell’ approximation (see e.g [14] or [34]). The coordinate $r$ is then replaced by $R$, except in the derivatives. For example $\frac{1}{r^{2}}\frac{\partial r^{2}w^{{\dagger}}}{\partial r}$ becomes $\frac{\partial w^{{\dagger}}}{\partial\mathsf{z}}$. With the above arguments, the Euler, mass conservation and incompressibility equations reduce to
$\left\{\begin{aligned} &\left[\frac{\partial}{\partial t}+\frac{u^{{\dagger}}}{R\cos\theta^{{\dagger}}}\frac{\partial}{\partial\varphi^{{\dagger}}}+\frac{v^{{\dagger}}}{R}\frac{\partial}{\partial\theta^{{\dagger}}}+w^{{\dagger}}\frac{\partial}{\partial\mathsf{z}}\right]\begin{pmatrix}u^{{\dagger}}\\ v^{{\dagger}}\\ w^{{\dagger}}\\ \end{pmatrix}+\\ &\qquad+2\Omega\begin{pmatrix}-v^{{\dagger}}\sin\varphi^{{\dagger}}\cos\theta^{{\dagger}}\\ u^{{\dagger}}\sin\varphi^{{\dagger}}\cos\theta^{{\dagger}}\\ 0\end{pmatrix}=-\frac{1}{\rho}\begin{pmatrix}\frac{1}{R\cos\theta^{{\dagger}}}\frac{\partial P}{\partial\varphi^{{\dagger}}}\\ \frac{1}{R}\frac{\partial P}{\partial\theta^{{\dagger}}}\\ \frac{\partial P}{\partial\mathsf{z}}\\ \end{pmatrix}-\begin{pmatrix}0\\ 0\\ g\\ \end{pmatrix},\\ &\,\frac{\partial\rho}{\partial t}+\frac{u^{{\dagger}}}{R\cos\theta^{{\dagger}}}\frac{\partial\rho}{\partial\varphi^{{\dagger}}}+\frac{v^{{\dagger}}}{R}\frac{\partial\rho}{\partial\theta^{{\dagger}}}+w^{{\dagger}}\frac{\partial\rho}{\partial\mathsf{z}}=0,\\ &\,\frac{1}{R\cos\theta^{{\dagger}}}\frac{\partial u^{{\dagger}}}{\partial\varphi^{{\dagger}}}+\frac{1}{R}\frac{\partial v^{{\dagger}}}{\partial\theta^{{\dagger}}}+\frac{\partial w^{{\dagger}}}{\partial\mathsf{z}}=0,\end{aligned}\right$. (2.23)
@@
Usually, the Coriolis parameter is defined by $f=2\Omega\sin\theta$, where $\theta\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ is the classical angle of latitude (see [45]), but, due to the relation $\sin\theta=\cos\theta^{{\dagger}}\sin\varphi^{{\dagger}}$ in (2.5) we can write